# Bulk Motions in the Black Hole Jet Sheath as a Candidate for the Comptonizing Corona

### v1

@@ 4 Discussion
Nonthermal X-ray emission is seen from AGN (Fabian et al., 2015) and X-ray binaries (Done et al., 2007). In previous papers (SB20, SSB21, SSB23, Gupta et al. 2024), we performed kinetic plasma (PIC) simulations of relativistic and trans-relativistic magnetic reconnection in $e^{\pm}$ and electron-ion plasmas. We found that the commonly-invoked scenario of thermal Comptonization of soft seed photons by hot electrons is unfeasible, since most of the electrons in the reconnection layer are cooled down to non-relativistic temperatures by the strong IC losses. However, the reconnection layer can still efficiently Comptonize photons. In fact, hard X-ray spectra in good agreement with the observations can be produced via Comptonization by the trans-relativistic bulk motions of the Compton-cooled plasmoid chain (à la Beloborodov, 2017). In this paper, we employ global GRRMHD simulations to explore which regions can produce the stochastic trans-relativistic bulk motions that seem required (based on our previous PIC studies) to act as a Comptonizing corona. Our main results are:
@@
    The sheath between the jet and the surrounding medium is an important site of energy dissipation, as revealed by the copious presence of reconnection layers and plasmoid chains. This region (which we have called ‘dissipative jet sheath’) is well identified by a combination of thresholds in plasma beta, $\beta>0.1$, and hot magnetization, $\sigma_{\rm h}>0.15$, i.e., more magnetized than the accretion disk yet less magnetized than the jet. The field lines that reconnect in this region are primarily radial. The time-averaged azimuthal field $B_{\rm out}$ in the dissipative jet sheath is comparable in strength to the reconnecting fields. While $B_{\rm out}$ does not always act as a guide field, we find that its polarity often stays constant across the current sheets located in the jet sheath (i.e., it can act as a guide field).
@@
    The radial electromagnetic power flowing in the jet sheath is twice larger than the electromagnetic power flowing in the jet core and is comparable to the overall accretion power. The electromagnetic power in the jet sheath decreases by $\sim 20\%$ from 2 to $10\,R_{\rm g}$. This may be able to satisfy the energetic requirements to explain the nonthermal X-rays from the hard state of black hole X-ray binaries (Remillard & McClintock, 2006; Done et al., 2007). We find that the polar angle of the jet sheath depends on distance as $\theta_{\rm pk}\simeq 0.3\pi(r/5\,R_{\rm g})^{-0.5}$, indicating a paraboloidal surface. Note that the X-ray-radio flux correlation ($F_{\rm radio}\propto F_{\rm X-ray}^{0.6}$) seen during the hard state, and the observed time lags were recently explained by invoking an outflowing, parabolic Comptonizing region (Kylafis et al., 2023; Kylafis & Reig, 2024).
@@
    The mean bulk motions in the jet sheath are trans-relativistic and dominated by the radial component, with time-averaged dimensionless 4-velocity $\langle\Gamma\beta_{\rm r}\rangle=1.2$. In the frame moving with the local mean velocity, the distribution of stochastic bulk motions resembles a Maxwellian with an effective bulk temperature of ${\cal O}$(100) keV, which is required to produce a Comptonized nonthermal X-ray spectrum in agreement with the observations.
@@
    The trans-relativistic radial bulk velocity of the jet sheath should have implications for the strength of the disk-reflected component (Beloborodov, 1999), the polarization angle (PA), and the PD. If the corona is launched away from the disk with speed $0.1\lesssim v/c\lesssim 0.8$—our measurements are in this range—Beloborodov (1998) found that the polarization angle would be parallel to the disk normal—this is the PA measured by IXPE for Cygnus X-1 (Krawczynski et al., 2022). Furthermore, IXPE observations revealed a large PD ($\sim 4\%$), which could be explained by an outflowing Comptonizing region with speed $v/c\gtrsim 0.4$ (akin to model (c) of Poutanen et al., 2023, and in the range of our measurements). Dexter & Begelman (2024) showed that the observed PA and PD of Cygnus X-1 could be explained by invoking a bulk outflow with $\Gamma\gtrsim 1.5$ along a conical sheath. Their requirement that the cone half-opening angle be $0.15\pi\lesssim\theta_{\rm cone}\lesssim 0.3\pi$ means that, in our model, most of the Comptonization must occur between 5 and 20 $R_{\rm g}$ (Fig. 7[c]).
@@
Several lines of observational evidence have pointed towards a link between the presence of a compact radio jet and the emission of nonthermal X-rays as seen during the low/hard states of X-ray binaries and AGN (Fender & Kuulkers, 2001; Gallo et al., 2003; Fender et al., 2004; Gallo et al., 2018; Méndez et al., 2022). Our candidate corona—the dissipative jet sheath—is naturally linked to the presence of a jet. As already discussed, our work is motivated by the radiative GRMHD simulations of Dexter et al. (2021); Liska et al. (2022), which showed the formation of an inner thicker disk (within $\sim 10\,R_{\rm g}$) with a strong poloidal field (akin to a MAD state) and a powerful jet, in case the flow is initially threaded with a dominant poloidal flux at large radii. It has been argued that the accretion disk might not be MAD during bright hard states based on the presence of Type-C quasi-periodic oscillations, which are thought to be associated to the precession of a thin disk (Ingram & Motta, 2019), and so not expected from a thicker, MAD disk (Fragile et al., 2023). We note, however, that Cygnus X-1 does not exhibit strong Type-C quasi-periodic oscillations during its hard state.
We conclude with a few caveats. The results presented in this paper have been obtained with two-dimensional simulations of accretion onto a highly spinning black hole in the MAD regime, i.e., with strong poloidal flux near the disk’s inner edge. In the future, it may be useful to investigate how our conclusions are influenced by: (1) a lower and possibly retrograde spin of the black hole; (2) three-dimensionality; (3) a dominant toroidal component of the magnetic field in the disk, which may weaken the jet; and (4) the self-consistent interaction between plasma and radiation. The latter is essential to properly determine the plasma composition, the resulting optical depth, and self-consistent Comptonized X-ray spectra.
@@ Acknowledgements
N.S. acknowledges the support from NASA (grant number 80NSSC22K0332), NASA FINESST (grant number 80NSSC22K1597), Columbia University Dean’s fellowship, and a grant from the Simons Foundation. B.R. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC). L.S. acknowledges support from the DoE Early Career Award DE-SC0023015. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to L.S. and B.R., and facilitated by Multimessenger Plasma Physics Center (MPPC), grant PHY-2206609 to L.S. JD is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51552.001A, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.
@@ Appendix B Jet-sheath efficiency
In Fig. 13, we check how the jet sheath efficiency $\eta_{\rm js}$ depends on different definitions for jet sheath and jet core, at our fiducial radius of $r=15\,R_{\rm g}$. The efficiency $\eta_{\rm js}$ presented in the main paper is defined as the ratio of polar-angle-integrated $\langle(E\times B)_{\rm r}\rangle$ in the jet sheath region to the jet core region. The solid black curve in Fig. 13 shows $(E\times B)_{\rm r}$ averaged during the quasi-steady state, as a function of $\theta$. One way we define the jet sheath is such that $\langle(E\times B)_{\rm r}\rangle[\theta]\geq 0.75\langle(E\times B)_{\rm r}\rangle[\theta_{\rm pk}]$, where $\theta_{\rm pk}$ is the polar angle where $\langle(E\times B)_{\rm r}\rangle$ peaks in the upper and lower hemispheres (blue patches); the jet core is the remaining green patch. For example, the $\theta$ ranges for the jet sheath region in the upper hemisphere going by this definition are: $0.25\pi\lesssim\theta\lesssim 0.39\pi$ at 5 $R_{\rm g}$, $0.17\pi\lesssim\theta\lesssim 0.27\pi$ at 10 $R_{\rm g}$, $0.15\pi\lesssim\theta\lesssim 0.22\pi$ at 15 $R_{\rm g}$, $0.13\pi\lesssim\theta\lesssim 0.18\pi$ at 20 $R_{\rm g}$, $0.11\pi\lesssim\theta\lesssim 0.16\pi$ at 25 $R_{\rm g}$, $0.10\pi\lesssim\theta\lesssim 0.14\pi$ at 30 $R_{\rm g}$, $0.09\pi\lesssim\theta\lesssim 0.13\pi$ at 35 $R_{\rm g}$, and $0.08\pi\lesssim\theta\lesssim 0.12\pi$ at 40 $R_{\rm g}$. This definition yields $\eta_{\rm js}\sim 2$, as discussed in the main paper. For other purposes, in the paper, we have defined the dissipative jet sheath as the region having $\sigma_{\rm h}>0.15,\beta>0.1$; regions satisfying these conditions only exist at certain times (which we shall call ‘active times’). We complement this by defining here the jet core as the region with $\sigma_{\rm h}>1,\beta<0.1$. The $\langle(E\times B)_{\rm r}\rangle$ averaged only during active times is shown by blue (jet sheath) and green (jet core) dotted curves. The corresponding jet sheath efficiency is $\eta_{\rm js}\sim 1.1$ (here, definitions of the jet sheath and jet core are based on our $\sigma_{\rm h}$ and $\beta$ thresholds). This would be $\eta_{\rm js}\sim 0.24$ if averaged over ‘all times’ during the quasi-steady state (dashed curves). We also find that the jet sheath efficiency increases by a factor of $\sim 2$ by considering the total radially flowing power $(E\times B)_{\rm r}(1+1/\sigma_{\rm h})$ instead of just the electromagnetic component $(E\times B)_{\rm r}$.
@@ Appendix C Electromagnetic dissipation along the jet sheath
Furthermore, we estimate the fraction of the radially-flowing electromagnetic power that is dissipated in the jet sheath: we find that the radially-flowing electromagnetic power in the jet sheath $\dot{E}^{\rm(EM)}(r)=2\pi\int\sqrt{g^{\rm rr}}(E\times B)_{\rm r}\sqrt{-g}{\rm d}\theta$ decreases with distance as $\propto r^{-0.2}$ between 2 and 10 $R_{\rm g}$ i.e., decreasing by $\sim 20\%$ in our candidate coronal region from $2-10\,R_{\rm g}$. We associate this fraction $\xi=0.2$ as the efficiency of converting the electromagnetic energy to plasma (and radiation), and find up to $\sim 10^{38}\,{\rm erg\,s^{-1}}$ to be dissipated between 2 and 10 $R_{\rm g}$ (for the parameters typical of Cygnus X-1; see §3.5). This is shown in Fig. 14; the dark blue solid curve is computed from the regions of the jet sheath defined by $\sigma_{\rm h}>0.15$ and $\beta>0.1$ (time-averaged during the active times, as mentioned above). As discussed in Qian et al. (2018), the increasing $\dot{E}^{\rm(EM)}(r)$ at larger radii $r\gtrsim 10\,R_{\rm g}$ could be due to the contribution of large-angled disk winds. This trend is robust regardless of whether the dissipative jet sheath is defined based on the $\sigma>0.15,\beta>0.1$ threshold or the $\langle(E\times B)_{\rm r}\rangle[\theta]\geq 0.75\langle(E\times B)_{\rm r}\rangle[\theta_{\rm pk}]$ condition.
Figure 12: Time-evolution of the radial Poynting flux as a function of $\theta$. Each panel denotes a different radius. All the shown quantities are the same as in Fig. 7[b].
Figure 13: Time-averaged radial Poynting flux $\langle(E\times B)_{\rm r}\rangle$ as a function of $\theta$ calculated at the fiducial radius of $r=15\,R_{\rm g}$. The blue and green dotted curves denote $\langle(E\times B)_{\rm r}\rangle$ computed only at the times when the $\sigma_{\rm h}>0.15,\beta>0.1$ criterion for jet sheath, and $\sigma_{\rm h}>1,\beta<0.1$ criterion for jet core are attained. The blue and green dashed curves denote $\langle(E\times B)_{\rm r}\rangle$ averaged over all times during the quasi-steady state. For dotted and dashed green lines, we consider only regions that satisfy $\sigma_{\rm h}>1,\beta<0.1$ (jet core). For dotted and dashed blue lines, we consider only regions that satisfy $\sigma_{\rm h}>0.15,\beta>0.1$ (jet sheath). The black curve denotes $\langle(E\times B)_{\rm r}\rangle$ averaged over all times in the whole volume; the blue and green patches define the jet sheath and jet core regions as done in the main text. The jet sheath efficiency $\eta_{\rm js}$ obtained for different definitions of jet sheath and jet core is also listed (see Appendix A for more details).

### v2

@@ 4 Discussion
Nonthermal X-ray emission is seen from active galactic nuclei (AGNs; Fabian et al., 2015) and X-ray binaries (Done et al., 2007). In previous papers (SB20, SSB21, SSB23, Gupta et al. 2024), we performed kinetic plasma (PIC) simulations of relativistic and trans-relativistic magnetic reconnection in $e^{\pm}$ and electron-ion plasmas. The simulations demonstrate that the popular scenario of hard X-ray production by thermal Comptonization is unfeasible for a corona heated by magnetic reconnection, since most of the electrons in the reconnection layer are kept at low temperatures by the strong Compton cooling. However, the reconnection layer can still produce the observed hard X-ray spectra via Comptonization by the stochastic bulk motions of the Compton-cooled plasmoid chain (Beloborodov, 2017). In this paper, we employ global GRRMHD simulations to explore which regions can produce the stochastic trans-relativistic plasmoid chain that could act as a Comptonizing corona. Our main results are as follows.
@@
    The sheath between the jet and the surrounding medium is an important site of energy dissipation, as revealed by the copious presence of reconnection layers and plasmoid chains. This region (which we have called the ‘dissipative jet sheath’) is well identified by a combination of thresholds in plasma beta, $\beta>0.1$, and hot magnetization, $\sigma_{\rm h}>0.15$, i.e., more magnetized than the accretion disk yet less magnetized than the jet.
@@
    The field lines that reconnect in this region are primarily radial. The time-averaged azimuthal field $B_{\rm out}$ in the dissipative jet sheath is comparable in strength to the reconnecting fields. While $B_{\rm out}$ does not always act as a guide field, we find that its polarity often stays constant across the current sheets located in the jet sheath (i.e., it can act as a guide field).
@@
    The radial electromagnetic power flowing in the jet sheath is twice as large as the electromagnetic power flowing in the jet core and is comparable to the overall accretion power. The electromagnetic power in the jet sheath decreases by $\sim 20\%$ from 2 to $10\,R_{\rm g}$. This may be able to satisfy the energetic requirements to explain the nonthermal X-rays from the hard state of black hole X-ray binaries (Remillard & McClintock, 2006; Done et al., 2007). We find that the polar angle of the jet sheath depends on distance as $\theta_{\rm pk}\simeq 0.3\pi(r/5\,R_{\rm g})^{-0.5}$, indicating a paraboloidal surface. Note that the X-ray-radio flux correlation ($F_{\rm radio}\propto F_{\rm X-ray}^{0.6}$) seen during the hard state, and the observed time lags were recently explained by invoking an outflowing, parabolic Comptonizing region (Kylafis et al., 2023; Kylafis & Reig, 2024).
@@
    The mean bulk motions in the jet sheath are trans-relativistic and dominated by the radial component, with time-averaged dimensionless 4-velocity $\langle\Gamma\beta_{\rm r}\rangle=1.2$. In the frame moving with the local mean velocity, the distribution of stochastic bulk motions resembles a Maxwellian with an effective bulk temperature of ${\cal O}$(100) keV, which is required to produce a Comptonized nonthermal X-ray spectrum in agreement with the observations. While the presence of trans-relativistic plasmoids from magnetic reconnection is a sufficient condition for stochastic bulk motion Comptonization, we emphasize that the stochasticity can also be potentially brought about by other means such as trans-relativistic turbulence (e.g., Grošelj et al., 2024) and/or Kelvin-Helmholtz-like vortices that are expected along the jet sheath (e.g., Wong et al., 2021; chow_23; Davelaar et al., 2023).
@@
    The trans-relativistic radial bulk velocity of the jet sheath should have implications for the strength of the disk-reflected component (Beloborodov, 1999), the polarization angle (PA), and the PD. If the corona is launched away from the disk with speed $0.1\lesssim v/c\lesssim 0.8$—our measurements are in this range—Beloborodov (1998) found that the PA would be parallel to the disk normal—this is the PA measured by IXPE for Cygnus X-1 (Krawczynski et al., 2022). Furthermore, IXPE observations revealed a large PD ($\sim 4\%$), which could be explained by an outflowing Comptonizing region with speed $v/c\gtrsim 0.4$ (akin to model (c) of Poutanen et al., 2023, and in the range of our measurements). Dexter & Begelman (2024) showed that the observed PA and PD of Cygnus X-1 could be explained by invoking a bulk outflow with $\Gamma\gtrsim 1.5$ along a conical sheath. Their requirement that the cone half-opening angle be $0.15\pi\lesssim\theta_{\rm cone}\lesssim 0.3\pi$ means that, in our model, most of the Comptonization must occur between 5 and 20 $R_{\rm g}$ (Fig. 7[c]).
  6. 6.
@@
Several lines of observational evidence have pointed toward a link between the presence of a compact radio jet and the emission of nonthermal X-rays as seen during the low/hard states of X-ray binaries and AGN (Fender & Kuulkers, 2001; Gallo et al., 2003; Fender et al., 2004; Gallo et al., 2018; Méndez et al., 2022). Our candidate corona—the dissipative jet sheath—is naturally linked to the presence of a jet. As already discussed, our work is motivated by the radiative GRMHD simulations of Dexter et al. (2021); Liska et al. (2022), which showed the formation of an inner thicker disk (within $\sim 10\,R_{\rm g}$) with a strong poloidal field (akin to a MAD state) and a powerful jet, in case the flow is initially threaded with a dominant poloidal flux at large radii. It has been argued that the accretion disk might not be MAD during bright hard states based on the presence of Type-C quasi-periodic oscillations, which are thought to be associated to the precession of a thin disk (Ingram & Motta, 2019), and so not expected from a thicker, MAD disk (Fragile et al., 2023). We note, however, that Cygnus X-1 does not exhibit strong Type-C quasi-periodic oscillations during its hard state.
We conclude by emphasizing limitations of the analysis presented in this paper. Our results are based on two-dimensional MHD simulations of accretion onto a highly spinning black hole in the MAD regime, i.e., with strong poloidal flux near the disk’s inner edge. We expect that the proposed Comptonization mechanism will also work in three-dimensional (3D) simulations, where both the kinetic drift-kink instability and the MHD flux-rope kink instability are enabled. This expectation is supported by existing local-box radiative 3D PIC simulations (Sironi & Beloborodov, 2020). In the future, it may be useful to investigate how our conclusions are influenced by: (1) three-dimensionality in global simulations; (2) a lower and possibly retrograde spin of the black hole; (3) a dominant toroidal component of the magnetic field in the disk, which may weaken the jet; (4) a spatially-dependent prescription for resistivity based on kinetic simulations of collisionless reconnection (Selvi et al., 2023; Bugli et al., 2024), which might yield a faster reconnection rate and may affect the properties of the plasmoid bulk motions in the jet sheath; and (5) the self-consistent interaction between plasma and radiation. The latter is essential to properly determine the plasma composition, the resulting optical depth, and self-consistent Comptonized X-ray spectra.
@@ Acknowledgements
N.S. acknowledges the support from NASA (grant No. 80NSSC22K0332), NASA FINESST (grant No. 80NSSC22K1597), Columbia University Dean’s fellowship, and a grant from the Simons Foundation. B.R. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC). L.S. acknowledges support from the DoE Early Career Award DE-SC0023015. This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470) to L.S. and B.R., and facilitated by Multimessenger Plasma Physics Center (MPPC), grant PHY-2206609 to L.S. J.D. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51552.001A, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. A.M.B. acknowledges support by NASA grants 80NSSC24K1229 and 21-ATP21-0056, NSF grant AST-2408199, and Simons Foundation grant No. 446228.
@@ Appendix B Jet-sheath efficiency
In Fig. 13, we check how the jet sheath efficiency $\eta_{\rm js}$ depends on different definitions for jet sheath and jet core, at our fiducial radius of $r=15\,R_{\rm g}$. The efficiency $\eta_{\rm js}$ presented in the main paper is defined as the ratio of polar-angle-integrated $\langle(E\times B)_{\rm r}\rangle$ in the jet sheath region to the jet core region. The solid black curve in Fig. 13 shows $(E\times B)_{\rm r}$ averaged during the quasi-steady state, as a function of $\theta$. One way we define the jet sheath is such that $\langle(E\times B)_{\rm r}\rangle[\theta]\geq 0.75\langle(E\times B)_{\rm r}\rangle[\theta_{\rm pk}]$, where $\theta_{\rm pk}$ is the polar angle where $\langle(E\times B)_{\rm r}\rangle$ peaks in the upper and lower hemispheres (blue patches); the jet core is the remaining green patch. For example, the ranges of $\theta$ for the jet sheath region in the upper hemisphere going by this definition are: $0.25\pi\lesssim\theta\lesssim 0.39\pi$ at 5 $R_{\rm g}$, $0.17\pi\lesssim\theta\lesssim 0.27\pi$ at 10 $R_{\rm g}$, $0.15\pi\lesssim\theta\lesssim 0.22\pi$ at 15 $R_{\rm g}$, $0.13\pi\lesssim\theta\lesssim 0.18\pi$ at 20 $R_{\rm g}$, $0.11\pi\lesssim\theta\lesssim 0.16\pi$ at 25 $R_{\rm g}$, $0.10\pi\lesssim\theta\lesssim 0.14\pi$ at 30 $R_{\rm g}$, $0.09\pi\lesssim\theta\lesssim 0.13\pi$ at 35 $R_{\rm g}$, and $0.08\pi\lesssim\theta\lesssim 0.12\pi$ at 40 $R_{\rm g}$. This definition yields $\eta_{\rm js}\sim 2$, as discussed in the main paper. For other purposes, in the paper, we have defined the dissipative jet sheath as the region having $\sigma_{\rm h}>0.15,\beta>0.1$; regions satisfying these conditions only exist at certain times (which we shall call ‘active times’). We complement this by defining here the jet core as the region with $\sigma_{\rm h}>1,\beta<0.1$. The $\langle(E\times B)_{\rm r}\rangle$ averaged only during active times is shown by blue (jet sheath) and green (jet core) dotted curves. The corresponding jet sheath efficiency is $\eta_{\rm js}\sim 1.1$ (here, definitions of the jet sheath and jet core are based on our $\sigma_{\rm h}$ and $\beta$ thresholds). This would be $\eta_{\rm js}\sim 0.24$ if averaged over ‘all times’ during the quasi-steady state (dashed curves). We also find that the jet sheath efficiency increases by a factor of $\sim 2$ by considering the total radially flowing power $(E\times B)_{\rm r}(1+1/\sigma_{\rm h})$ instead of just the electromagnetic component $(E\times B)_{\rm r}$.
@@ Appendix C Electromagnetic dissipation along the jet sheath
Furthermore, we estimate the fraction of the radially-flowing electromagnetic power that is dissipated in the jet sheath: we find that the radially-flowing electromagnetic power in the jet sheath $\dot{E}^{\rm(EM)}(r)=2\pi\int\sqrt{g^{\rm rr}}(E\times B)_{\rm r}\sqrt{-g}{\rm d}\theta$ decreases with distance as $\propto r^{-0.2}$ between 2 and 10 $R_{\rm g}$ i.e., it decreases by $\sim 20\%$ in our candidate coronal region from $2-10\,R_{\rm g}$. We associate this fraction $\xi=0.2$ with the efficiency of converting the electromagnetic energy to plasma (and radiation), and find up to $\sim 10^{38}\,{\rm erg\,s^{-1}}$ to be dissipated between 2 and 10 $R_{\rm g}$ (for the parameters typical of Cygnus X-1; see §3.5). This is shown in Fig. 14; the dark blue solid curve is computed from the regions of the jet sheath defined by $\sigma_{\rm h}>0.15$ and $\beta>0.1$ (time-averaged during the active times, as mentioned above). As discussed by Qian et al. (2018), the increasing $\dot{E}^{\rm(EM)}(r)$ at larger radii $r\gtrsim 10\,R_{\rm g}$ could be due to the contribution of large-angled disk winds. This trend is robust regardless of whether the dissipative jet sheath is defined based on the $\sigma>0.15,\beta>0.1$ threshold or the $\langle(E\times B)_{\rm r}\rangle[\theta]\geq 0.75\langle(E\times B)_{\rm r}\rangle[\theta_{\rm pk}]$ condition.
Figure 12: Time-evolution of the radial Poynting flux as a function of $\theta$. Each panel denotes a different radius. All the quantities shown are the same as in Fig. 7[b].
Figure 13: Time-averaged radial Poynting flux $\langle(E\times B)_{\rm r}\rangle$ as a function of $\theta$ calculated at the fiducial radius of $r=15\,R_{\rm g}$. The blue and green dotted curves denote $\langle(E\times B)_{\rm r}\rangle$ computed only at the times when the criteria $\sigma_{\rm h}>0.15,\beta>0.1$ for jet sheath, and $\sigma_{\rm h}>1,\beta<0.1$ criterion for jet core are attained. The blue and green dashed curves denote $\langle(E\times B)_{\rm r}\rangle$ averaged over all times during the quasi-steady state. For dotted and dashed green lines, we consider only regions that satisfy $\sigma_{\rm h}>1,\beta<0.1$ (jet core). For dotted and dashed blue lines, we consider only regions that satisfy $\sigma_{\rm h}>0.15,\beta>0.1$ (jet sheath). The black curve denotes $\langle(E\times B)_{\rm r}\rangle$ averaged over all times in the whole volume; the blue and green patches define the jet sheath and jet core regions as done in the main text. The jet sheath efficiency $\eta_{\rm js}$ obtained for different definitions of jet sheath and jet core is also listed (see Appendix B for more details).